# ZMP-HH-25/5

### v1

@@ 5.2 Comparison of Worldsheet Theories
$(\mathbf{b}^{0},\mathbf{b}^{1},\mathbf{b}^{2},\mathbf{b}^{3},u)\quad\longleftrightarrow\quad(\widehat{\mathbf{b}}^{0},\widehat{\mathbf{b}}^{1},\widehat{\mathbf{b}}^{2},\widehat{\mathbf{b}}^{3},\widehat{u})\,.$ (5.8)
@@
$\widehat{\mathbf{b}}^{0},\widehat{\mathbf{b}}^{3}$ $\widehat{\Lambda}_{\mathrm{trans}}$ interacting
@@
$\mathbf{b}^{0}$ reduction $B_{2}$ interacting
@@
$n_{\mathrm{V}}^{\rm IIA}$ =$\stackrel{{\scriptstyle!}}{{=}}n_{\rm V}^{\rm pert}$ (5.12)
@@
At the same time, the number of vector multiplets in Type IIA compactified on $\widehat{V}$ is given by $h^{1,1}(\widehat{V})$. Hence, we recover the result that there are no additional fibral divisors localised over points in the base of the K3-fibred threefold $\widehat{V}$.
@@
$n_{\mathrm{V}}^{\rm IIB}$ =$\stackrel{{\scriptstyle!}}{{=}}n_{\rm V}^{\rm pert}$ (5.13)
@@
Let us now return to the conditions under which $\mathtt{S}_{Z}$ and $\widehat{\mathtt{S}}_{\widehat{Z}}$ give rise to the same perturbative heterotic string. First, the perturbative gauge groups must agree. For the perturbative heterotic string by (5.12) and (5.13) this is the case if $h^{2,1}(V)=h^{1,1}(\widehat{V})$. Next, we require that the dimension of the moduli space of the heterotic gauge bundle is the same. This is a necessary condition for the gauge bundles of $\mathtt{S}_{Z}$ and $\widehat{\mathtt{S}}_{\widehat{Z}}$ (and thus the interactions of the worldsheet theory) to agree. Since we are focusing on the perturbative heterotic string for which all massless hypermultiplets are perturbative string excitations, this imposes the condition $h^{2,1}(\widehat{V})=h^{1,1}(V)$. Together, both conditions are satisfied if $\widehat{V}$ and $V$ are mirror to each other.
To summarise, for the worldsheet theory of the strings $\mathtt{S}_{Z}$ and $\widehat{\mathtt{S}}_{\widehat{Z}}$ to be the same, including interactions, the manifolds $V$ and $\widehat{V}$ on which Type IIB and IIA string theory are respectively compactified, have to be mirror to each other. Conversely, consider a Calabi–Yau threefold $V$ that allows for a Tyurin degeneration giving rise to a string $\mathtt{S}_{Z}$ if we compactify Type IIB string theory on it. Then the mirror $\widehat{V}$ has to be K3-fibred with generic fibre $\widehat{Z}$ such that Type IIA compactified on $\widehat{V}$ has an emergent string with the same worldsheet theory as $\mathtt{S}_{Z}$. This is consistent with the expectation of [55]: A Calabi–Yau threefold $V$ which allows for a Tyurin degeneration of the form $V\to X_{1}\cup_{Z}X_{2}$ such that $X_{1}$ and $X_{2}$ induce a lattice polarization on $Z$ of rank $\rho+1$ is mirror dual to a threefold $\widehat{V}$ allowing for a fibration by a K3-surface $\widehat{Z}$ with polarization lattice of rank $19-\rho$.
@@ 5.3 Direct Application of Mirror Symmetry to NS5-Brane String
Consider again the K3-fibred Calabi–Yau threefold $\widehat{V}$ with generic fibre $\widehat{Z}$, as defined in (5.1). To identify the SYZ special Lagrangian $T^{3}$-fibre, we first notice that $\widehat{Z}$ itself allows for a fibration by a calibrated curve $\widehat{\mathcal{E}}$ that is topologically a $T^{2}$ and satisfies $J_{\widehat{Z}}|_{\widehat{\mathcal{E}}}=0$. Since $\widehat{\mathcal{E}}$ is not a holomorphic curve, it is not defined globally on $\widehat{V}$, but is subject to monodromies around the points in ^$\widehat{\mathbb{P}}^{1}_{b}$ over which $\widehat{Z}$ degenerates. Still, if there exists a non-contractible path $\widehat{\gamma}\subset\widehat{\mathbb{P}}^{1}_{b}$ along which $\widehat{\mathcal{E}}$ does not undergo a monodromy, the restriction of the K3 fibration to $\widehat{\gamma}$ gives a 3-cycle. However, since $h^{1}(\widehat{\mathbb{P}}^{1}_{b})=0$ no such $\widehat{\gamma}$ exists on ^$\widehat{\mathbb{P}}^{1}_{b}$, but there can exist multi-covers of ^$\widehat{\mathbb{P}}^{1}_{b}$ that are branched over the degeneration points of $\widehat{Z}$ in ^$\widehat{\mathbb{P}}^{1}_{b}$ with a different topology. In the following, we discuss one such instance.
@@
See Figure -955(a) for a geometric depiction. To avoid confusion, let us stress that for $\widehat{V}$ we are not interested in the physics of the Tyurin degeneration since we consider the vector multiplet moduli space of Type IIA compactified on $\widehat{V}$. In the sequel, we merely need to assume that there exists such a Tyurin degeneration for $\widehat{V}$. For a large class of examples with this property, see [56]. Notice that for the Tyurin degeneration depicted in -955(a) this requires that the degeneration points of the K3-fibration $\widehat{V}$ can be separated into two sets such that the net monodromy of the fibre associated with each subset is trivial.
Figure -955: (a) The Tyurin degeneration (5.14) in which the base of the K3-fibration splits into two components ^$\widehat{\mathbb{P}}^{1}_{1}$ and ^$\widehat{\mathbb{P}}^{1}_{2}$ intersecting over a single point. (b) The same situation as in Figure (a), but with the intersection point removed, giving rise to two asymptotically cylindrical geometries the 1-cycles $\widehat{\gamma}_{1}$ and $\widehat{\gamma}_{2}$ at the boundary.
Since $\widehat{X}_{1}$ is a half Calabi–Yau threefold, we can remove a point in the base, $\widehat{z}_{1}\in\widehat{\mathbb{P}}^{1}_{1}$, of $\widehat{X}_{1}$ over which the fibre $\widehat{Z}_{1}$ is smooth and glue in one boundary of a cylinder at $\widehat{z}_{1}$. Let us denote the boundary of the cylinder by $\widehat{\gamma}_{1}$. Similarly, we can replace a point in the base of $\widehat{X}_{2}$ by a cylinder with boundary $\widehat{\gamma}_{2}$. Since both $\widehat{X}_{1}$ and $\widehat{X}_{2}$ thus have asymptotically cylindrical regions, the monodromy generated by transporting $\widehat{Z}_{1/2}$ around $\widehat{\gamma}_{1/2}$ is trivial. See also Figure -955(b). We obtain $\widehat{V}_{0}$ by gluing the asymptotic regions of $\widehat{X}_{1}$ and $\widehat{X}_{2}$. Let $\widehat{i}_{k}:\widehat{X}_{k}\hookrightarrow\widehat{V}_{0}$, $k=1,2$ be the embeddings of $\widehat{X}_{k}$ in $\widehat{V}_{0}$. On $\widehat{V}_{0}$ we can then consider the one-cycle
@@
where the minus sign takes care of the orientation of $\widehat{\gamma}_{1,2}$ such that the cycle $\widehat{\gamma}$ is nontrivial. Since ^$\widehat{\mathbb{P}}^{1}_{b}$ has no one-cycles, $\widehat{\gamma}$ cannot be viewed as a one-cycle on ^$\widehat{\mathbb{P}}^{1}_{b}$ identified as a section of the K3-fibration. Instead, in the simplest case, $\widehat{\gamma}$ is a one-cycle on a bi-section $\widehat{B}$ of the K3 fibration that is a double cover of ^$\widehat{\mathbb{P}}^{1}_{b}$. Since this bi-section has $h^{1}(\widehat{B})>0$, it is a higher-genus curve. For this to be possible, the bi-section must be irreducible, which can occur due to branching over the points in ^$\widehat{\mathbb{P}}^{1}_{b}$ over which $\widehat{Z}$ degenerates. The simplest case would be that the bi-section is topologically a $T^{2}$ with one of the one-cycles of the $T^{2}$ given by the path $\widehat{\gamma}$. For definiteness, we will make this additional assumption from now on. Since the monodromy around $\widehat{\gamma}_{1,2}$ is trivial, also the monodromy around $\widehat{\gamma}$ is trivial such that ^×$\widehat{\mathcal{E}}\times\widehat{\gamma}$ gives a well-defined 3-cycle of $T^{3}$ topology. The situation is depicted schematically in Figure -954(a).
Figure -954: The geometry of $\widehat{V}$ (a) and its mirror $V$ (b). Indicated is the one-cycle $\widehat{\gamma}$ that lives on the double cover of the base ^$\widehat{\mathbb{P}}^{1}_{b}$ of $\widehat{V}$ and its mirror $\gamma$ on the genus one-fibre of $V$. Furthermore the K3-fibre $\widehat{Z}$ and its mirror $Z$ realised as a bi-section of the genus-one fibration of $V$ are indicated in green.
@@
Consider now the large base limit $\text{vol}(\widehat{\mathbb{P}}^{1}_{b})\to\infty$ for the threefold $\widehat{V}$. In this limit $\widehat{\gamma}\in H_{1}(\widehat{B})$ grows to infinite length. Hence, in the mirror dual of the large base limit, the dual cycle $\gamma\in H_{1}(B)$ pinches. Let us further assume that the normal bundle of $B$ inside $V$ is trivial, $\mathcal{N}_{B|V}=\mathcal{O}(0)\oplus\mathcal{O}(0)$. Then, apart from the fibration with generic fiber $Z$, the threefold $V$ allows for a second, incompactible genus-one fibration with generic fiber $B$. An example of such a geometry can be found in [90]. In this case, the mirror dual of the large base limit for $\widehat{V}$ corresponds to the limit in which the generic fibre of the genus-one fibration of $V$ degenerates.
@@
Given the split of mirror symmetry for $\widehat{V}$ into mirror symmetry for $\widehat{Z}$ and a T-duality along $\widehat{\gamma}$, the NS5-brane wrapping $\widehat{Z}$ is not affected by the application of K3-mirror symmetry to $\widehat{Z}$. Instead, T-duality along $\widehat{\gamma}$ turns the NS5-brane into a KK-monopole. Four directions of the worldvolume theory of the monopole are hence identified with $Z$. Moreover, the cycle $\gamma$ dual to $\widehat{\gamma}$ takes over the role of the KK-circle and shrinks at the location of the monopole. The monopole geometry is illustrated in Figure -954(b). At the centre of the monopole, the genus-one fibre of $V$ is hence degenerate. Since $\widehat{\gamma}$ is a one-cycle on the double-cover of ^$\widehat{\mathbb{P}}^{1}_{b}$, the worldvolume of the monopole is a double-cover of the base $B_{2}$ of $V$ viewed as a genus-one fibration. Moreover, at the location of the KK-monopole the discriminant of the genus-one curve $B$ has a double zero. Altogether, we conclude that at the centre of the monopole we obtain a degenerate Calabi–Yau threefold $V_{0}$ that is a genus-one fibration with a bi-section identified as $Z$. The degeneration amounts to an $I_{2}$ singularity in the generic genus-one fibre. In F-theory language, the centre of the KK-monopole hence realises a Sen-limit [91]. [^40] See [12, 13, 16, 15] for the description of type II limits of elliptic K3 and Calabi-Yau threefolds as Sen-limits. After resolving the $I_{2}$ singularity, the fibre $V_{0}$ over the centre of the monopole splits into the union of two Fano threefolds
@@ 6.1 Type II Degenerations with Abelian Surfaces
To determine the spectrum on the worldsheet of the EFT string realising the type II degeneration of the form (6.2), we proceed as in Section 4 and study the modes that localise on $V_{0}$. First, the zero modes $(\mathbf{z}_{0},\mathbf{b}^{0},\widetilde{\mathbf{b}}^{0},\Phi)$ corresponding to the directions transverse to $A$ in $\mathcal{V}$ as well as those obtained from $B_{6}$ and $C_{6}$ are universally present and hence give rise to five left- and five right-moving scalars on the worldsheet. Additional zero modes arise from reducing $C_{4}$ over elements in $H^{2}(A)$. For $A$ an Abelian surface, $H^{2}(A,\mathbb{Z})$ can be identified with the lattice $\Gamma_{3,3}$. In total this gives rise to three left- and three right-moving scalars, making a total of eight left- and eight right-moving scalars. As before the string solution is a BPS solution leading to an $\mathcal{N}=(0,4)$ superalgebra that is manifest on the string. Therefore, as in Section 4, the right-moving scalar fields are accompanied by right-moving fermions to form full $\mathcal{N}=(0,4)$ multiplets. However, unlike for the case discussed in Section 4, for the degeneration (6.2) there are additional left-moving fermions arising from reducing the 10d gravitinos over one-cycles localised on $V_{0}$. The number of the left-moving fermions arising in this way is counted by $b_{1}(A)$, which for an Abelian surface gives $b_{1}(A)=4$ fermionic zero modes per gravitino, hence eight left-moving fermions in total. The spectrum of the EFT string associated with the degeneration (6.2) thus gives rise to a 2d worldsheet theory with central charges

### v2

@@ 5.2 Comparison of Worldsheet Theories
[…]
@@
$\widehat{\mathbf{b}}^{0},\widehat{\mathbf{c}}^{0}$ $\widehat{\Lambda}_{\mathrm{trans}}$ interacting
@@
$\mathbf{b}^{0},\mathbf{c}^{0}$ reduction $B_{2},C_{2}$ interacting
@@
$n_{\mathrm{V}}^{\rm IIA}$ $\mathrel{{\mathop{=}\limits^{!}}}n_{\rm V}^{\rm pert}$ (5.12)
@@
[…]
@@
$n_{\mathrm{V}}^{\rm IIB}$ $\mathrel{{\mathop{=}\limits^{!}}}n_{\rm V}^{\rm pert}$ (5.13)
@@
[…]
To summarise, for the worldsheet theory of the strings $\mathtt{S}_{Z}$ and $\widehat{\mathtt{S}}_{\widehat{Z}}$ to be the same, including interactions, the manifolds $V$ and $\widehat{V}$ on which Type IIB and, respectively, IIA string theory are compactified have to be mirror to each other. Conversely, consider a Calabi–Yau threefold $V$ that allows for a Tyurin degeneration giving rise to a string $\mathtt{S}_{Z}$ upon compactifying Type IIB string theory on it. Then the mirror $\widehat{V}$ has to be K3-fibred with generic fibre $\widehat{Z}$ such that Type IIA compactified on $\widehat{V}$ has an emergent string with the same worldsheet theory as $\mathtt{S}_{Z}$. This is consistent with the expectation of [55]: A Calabi–Yau threefold $V$ which allows for a Tyurin degeneration of the form $V\to X_{1}\cup_{Z}X_{2}$ such that $X_{1}$ and $X_{2}$ induce a lattice polarization on $Z$ of rank $\rho+1$ is mirror dual to a threefold $\widehat{V}$ allowing for a fibration by a K3-surface $\widehat{Z}$ with polarization lattice of rank $19-\rho$.
@@ 5.3 Direct Application of Mirror Symmetry to NS5-Brane String
Consider again the K3-fibred Calabi–Yau threefold $\widehat{V}$ with generic fibre $\widehat{Z}$, as defined in (5.1). To identify the SYZ special Lagrangian $T^{3}$-fibre, we first notice that $\widehat{Z}$ itself allows for a fibration by a calibrated curve $\widehat{\mathcal{E}}$ that is topologically a $T^{2}$ and satisfies $J_{\widehat{Z}}|_{\widehat{\mathcal{E}}}=0$. Since $\widehat{\mathcal{E}}$ is not a holomorphic curve, it is not defined globally on $\widehat{V}$, but is subject to monodromies around the points in $\widehat{\mathbb{P}}^{1}_{b}$ over which $\widehat{Z}$ degenerates. Still, if there exists a non-contractible path $\widehat{\gamma}\subset\widehat{\mathbb{P}}^{1}_{b}$ along which $\widehat{\mathcal{E}}$ does not undergo a monodromy, the restriction of the K3 fibration to $\widehat{\gamma}$ gives a 3-cycle. However, since $h^{1}(\widehat{\mathbb{P}}^{1}_{b})=0$ no such $\widehat{\gamma}$ exists on $\widehat{\mathbb{P}}^{1}_{b}$, but there can exist multi-covers of $\widehat{\mathbb{P}}^{1}_{b}$ that are branched over the degeneration points of $\widehat{Z}$ in $\widehat{\mathbb{P}}^{1}_{b}$ with a different topology. In the following, we discuss one such instance.
@@
See Figure 5(a) for a geometric depiction. To avoid confusion, let us stress that for $\widehat{V}$ we are not interested in the physics of the Tyurin degeneration since we consider the vector multiplet moduli space of Type IIA compactified on $\widehat{V}$. In the sequel, we merely need to assume that there exists such a Tyurin degeneration for $\widehat{V}$. For a large class of examples with this property, see [57, 56]. Notice that for the Tyurin degeneration depicted in 5(a) this requires that the degeneration points of the K3-fibration $\widehat{V}$ can be separated into two sets such that the net monodromy of the fibre associated with each subset is trivial.
Figure 5: (a) The Tyurin degeneration (5.14) in which the base of the K3-fibration splits into two components $\widehat{\mathbb{P}}^{1}_{1}$ and $\widehat{\mathbb{P}}^{1}_{2}$ intersecting over a single point. (b) The same situation as in Figure (a), but with the intersection point removed, giving rise to two asymptotically cylindrical geometries the 1-cycles $\widehat{\gamma}_{1}$ and $\widehat{\gamma}_{2}$ at the boundary.
Since $\widehat{X}_{1}$ is a half Calabi–Yau threefold, we can remove a point in the base, $\widehat{z}_{1}\in\widehat{\mathbb{P}}^{1}_{1}$, of $\widehat{X}_{1}$ over which the fibre $\widehat{Z}_{1}$ is smooth and glue in one boundary of a cylinder at $\widehat{z}_{1}$. Let us denote the boundary of the cylinder by $\widehat{\gamma}_{1}$. Similarly, we can replace a point in the base of $\widehat{X}_{2}$ by a cylinder with boundary $\widehat{\gamma}_{2}$. Since both $\widehat{X}_{1}$ and $\widehat{X}_{2}$ thus have asymptotically cylindrical regions, the monodromy generated by transporting $\widehat{Z}_{1/2}$ around $\widehat{\gamma}_{1/2}$ is trivial. See also Figure 5(b). We obtain $\widehat{V}_{0}$ by gluing the asymptotic regions of $\widehat{X}_{1}$ and $\widehat{X}_{2}$. Let $\widehat{i}_{k}:\widehat{X}_{k}\hookrightarrow\widehat{V}_{0}$, $k=1,2$ be the embeddings of $\widehat{X}_{k}$ in $\widehat{V}_{0}$. On $\widehat{V}_{0}$ we can then consider the one-cycle
@@
where the minus sign takes care of the orientation of $\widehat{\gamma}_{1,2}$ such that the cycle $\widehat{\gamma}$ is nontrivial. Since $\widehat{\mathbb{P}}^{1}_{b}$ has no one-cycles, $\widehat{\gamma}$ cannot be viewed as a one-cycle on $\widehat{\mathbb{P}}^{1}_{b}$ identified as a section of the K3-fibration. Instead, in the simplest case, $\widehat{\gamma}$ is a one-cycle on a bi-section $\widehat{B}$ of the K3 fibration that is a double cover of $\widehat{\mathbb{P}}^{1}_{b}$. Since this bi-section has $h^{1}(\widehat{B})>0$, it is a higher-genus curve. For this to be possible, the bi-section must be irreducible, which can occur due to branching over the points in $\widehat{\mathbb{P}}^{1}_{b}$ over which $\widehat{Z}$ degenerates. The simplest case would be that the bi-section is topologically a $T^{2}$ with one of the one-cycles of the $T^{2}$ given by the path $\widehat{\gamma}$. For definiteness, we will make this additional assumption from now on. Since the monodromy around $\widehat{\gamma}_{1,2}$ is trivial, also the monodromy around $\widehat{\gamma}$ is trivial such that $\widehat{\mathcal{E}}\times\widehat{\gamma}$ gives a well-defined 3-cycle of $T^{3}$ topology. The situation is depicted schematically in Figure 6(a).
Figure 6: The geometry of $\widehat{V}$ (a) and its mirror $V$ (b). Indicated is the one-cycle $\widehat{\gamma}$ that lives on the double cover of the base $\widehat{\mathbb{P}}^{1}_{b}$ of $\widehat{V}$ and its mirror $\gamma$ on the genus one-fibre of $V$. Furthermore the K3-fibre $\widehat{Z}$ and its mirror $Z$ realised as a bi-section of the genus-one fibration of $V$ are indicated in green.
@@
[…]
@@
Given the split of mirror symmetry for $\widehat{V}$ into mirror symmetry for $\widehat{Z}$ and a T-duality along $\widehat{\gamma}$, the NS5-brane wrapping $\widehat{Z}$ is not affected by the application of K3-mirror symmetry to $\widehat{Z}$. Instead, T-duality along $\widehat{\gamma}$ turns the NS5-brane into a KK-monopole. Four directions of the worldvolume theory of the monopole are hence identified with $Z$. Moreover, the cycle $\gamma$ dual to $\widehat{\gamma}$ takes over the role of the KK-circle and shrinks at the location of the monopole. The monopole geometry is illustrated in Figure 6(b). At the centre of the monopole, the genus-one fibre of $V$ is hence degenerate. Since $\widehat{\gamma}$ is a one-cycle on the double-cover of $\widehat{\mathbb{P}}^{1}_{b}$, the worldvolume of the monopole is a double-cover of the base $B_{2}$ of $V$ viewed as a genus-one fibration. Moreover, at the location of the KK-monopole the discriminant of the genus-one curve $B$ has a double zero. Altogether, we conclude that at the centre of the monopole we obtain a degenerate Calabi–Yau threefold $V_{0}$ that is a genus-one fibration with a bi-section identified as $Z$. The degeneration amounts to an $I_{2}$ singularity in the generic genus-one fibre. In F-theory language, the centre of the KK-monopole hence realises a Sen-limit [93]. [^41] See [12, 13, 16, 15] for the description of type II limits of elliptic K3 and Calabi-Yau threefolds as Sen-limits. After resolving the $I_{2}$ singularity, the fibre $V_{0}$ over the centre of the monopole splits into the union of two Fano threefolds
@@ 6.1 Type II Degenerations with Abelian Surfaces
To determine the spectrum on the worldsheet of the EFT string realising the type II degeneration of the form (6.2), we proceed as in Section 4 and study the modes that localise on $V_{0}$. First, the zero modes $(\mathbf{z}_{0},\mathbf{b}^{0},\mathbf{c}^{0},\Phi)$ corresponding to the directions transverse to $A$ in $\mathcal{V}$ as well as those obtained from $B_{6}$ and $C_{6}$ are universally present and hence give rise to five left- and five right-moving scalars on the worldsheet. Additional zero modes arise from reducing $C_{4}$ over elements in $H^{2}(A)$. For $A$ an Abelian surface, $H^{2}(A,\mathbb{Z})$ can be identified with the lattice $\Gamma_{3,3}$. In total this gives rise to three left- and three right-moving scalars, making a total of eight left- and eight right-moving scalars. As before the string solution is a BPS solution leading to an $\mathcal{N}=(0,4)$ superalgebra that is manifest on the string. Therefore, as in Section 4, the right-moving scalar fields are accompanied by right-moving fermions to form full $\mathcal{N}=(0,4)$ multiplets. However, unlike for the case discussed in Section 4, for the degeneration (6.2) there are additional left-moving fermions arising from reducing the 10d gravitinos over one-cycles localised on $V_{0}$. The number of the left-moving fermions arising in this way is counted by $b_{1}(A)$, which for an Abelian surface gives $b_{1}(A)=4$ fermionic zero modes per gravitino, hence eight left-moving fermions in total. The spectrum of the EFT string associated with the degeneration (6.2) thus gives rise to a 2d worldsheet theory with central charges